You roll a pair of dice five times. Find the following probability of you getting at LEAST one sum of 7.
I have trouble with the 'at least' and 'at most' probability problems.
is it, 5C1(6/36)^1 (1-30/36)^4 ?
I'm also unsure when to subtract the probability from one or when to not.
I also have to find the probability of getting at most one sum of 7.

Answer by sudhanshu_kmr(1152)

here at least means : in five times, at least one time the sum should be 7, it may be more than one times even all the five times.

probability of getting sum of 7 in a pair of dice = 6/36 = 1/6
probability of not getting sum of 7 = 5/6

probability of getting sum of 7 not a single time P(A) = 5C5(5/6)^5(1/6)^0
= 3125/7776

probability of getting sum of 7 at least one time P'(A)= 1 - P(A)
= 1 - 3125/7776
= 4651/7776

you are welcome to clear any doubt..